OpenStudy (anonymous):

(solve the equation for exact values of x, 0 is less than or equal to x which is less than 2 pi) 2tan^2(x)-3secx=0

OpenStudy (anonymous):

Convert tanx and secx interms of sinx and cosx

OpenStudy (anonymous):

Can u do that?

OpenStudy (anonymous):

I converted 2tan^2(x) into 2(sec^2(x)-1) so I got (2sec^2(x)-2)-3secx=0 then broke apart 2sec^2(x) to (1/cosx*2/cosx)-2-(3/cosx)=0 then multiplied by cos

OpenStudy (anonymous):

2tan^2(x)-3secx=0 2(sec^2x-1)-3secx=0 2sec^2x-3secx-2=0

OpenStudy (anonymous):

I did ad 3/cosx to both siwhich left me with 1/cosx*2/cosx-2=3/cosx then I multiplied by cos x 1*2-2cos=3 subtract 2 from both sides = -2cos=1 then divide by -2= cos=-1/2

OpenStudy (anonymous):

Now, factor it

OpenStudy (anonymous):

is cos = -1/2